# Investigating the Electroweak Phase Transition with a Real Scalar Singlet at a Muon Collider

### v1

@@ 4.2 $4l$ Final States
The four-lepton final state originates from $S\rightarrow ZZ$, and its dominant background is $\mu^{+}\mu^{-}\rightarrow 2l^{+}2l^{-}\mu^{+}\mu^{-}$ and $\mu^{+}\mu^{-}\rightarrow 2l^{+}2l^{-}\nu_{\mu}\bar{\nu}_{\mu}$. The signal leptons are expected to have higher transverse momentum compared to the background ones, as they originate from $Z$-boson decays, which suggests a constraint on the transverse momentum in terms of $Z$-boson mass as $\sim P_{T}(l^{\pm})\geq\frac{2}{5}M_{z}$. This assumption was verified by comparing the signal-background transverse momentum distributions for the charged leptons, which showed signal domination in the region $\sim 50\,\mathrm{GeV}\leq P_{T}(l^{\pm})\leq 400\,\mathrm{GeV}$. Consequently, we identified any oppositely-charged same-flavour leptons whose $P_{T}(l^{\pm})\geq\frac{2}{5}M_{z}$, and their invariant mass peaked around the $Z$-boson mass, $0.8M_{Z}\leq M(l^{+}l^{-})\leq 1.2M_{Z}$ as being valid pairs resulting from $Z-$boson decay. This works as a strong signal-background separator because the leptons produced from the background generated in MG5_aMC@NLO are sourced from different mediators, as discussed in the previous section. This is in contrast to the signal case where they are mainly produced from $Z$-bosons, $S\rightarrow ZZ$. Then, the surviving pairs were further constrained by requiring the invariant mass of the two pairs to be in the vicinity of the scalar mass, $M_{s}-50\,\mathrm{GeV}\leq M(l^{+}l^{-}l^{+}l^{-})\leq M_{s}+50\,\mathrm{GeV}$, which further suppressed the background events, as they were not resonantly produced in this range.
@@
Finally, the resulting sensitivity plot for a luminosity of $10^{4}\,\mathrm{fb}^{-1}$ is shwon in Fig. 9, at $68\%$ and $95\%$ C.L.
@@ 4.3 $4q$ Final States
The main background is similar to those discussed in the previous channel, with $l\rightarrow q$ now. In principle, the jets originating from ($ZZ,\,\,W^{+}W^{-},\,\,hh$) decays will have almost identical kinematics due to the approximate symmetry in the $Z,W^{\pm},h$ masses especially for $M_{S}\gg M_{Z},M_{W^{\pm}},M_{h}$. Nevertheless, the differences in the branching ratios of Higgs boson and weak gauge boson decays into quarks will lead to a deviation in the kinematical behavior of the two cases, especially for the transverse momentum and invariant mass distributions, as can be seen in Fig. 10 (A,C). These differences can be further investigated by tracking the jet’s origin, and this can play a vital role in exploring the $\mathbb{Z}_{2}$ nature of xSM. Hence, this specific channel, which has not been explored sufficiently to date, to the best of our knowledge, can further constrain the parameter space of FOEWPT. This proves as an additional advantage of a muon collider, in which such channel could be explored with much less effort than at hadron colliders due to the lower QCD backgrounds. A detailed study of these differences left for future work. Jets were clustered using FastJet (v3.3.2) Cacciari:2011ma , where the anti-kT algorithm Catani:1993hr ; Ellis:1993tq ; Cacciari:2008gp with a radius parameter $R=0.4$ was chosen to be the default jet clustering algorithm. The calculation of the invariant mass for this channel is more challenging than in the previous one, mainly due to the high tendency of quarks to radiate via QCD. This means that we can not simply calculate the invariant mass as previously using ${M^{2}(4j)=\big{(}\sum_{i=1}^{4}p(J_{i})\big{)}^{2}}$. Instead, for each event, we defined ${P=\sum_{i=1}^{n}p(J_{i})}$, that sums all the jets to a single four vector, from which we construct $M^{2}(nj)=P^{2}$. This, still, will not add up to the parent $S$ particle mass, as some of the energy will have escaped in the form of radiation, but this could be neglected as a similar effect will be also found in background events. The main effect of this final-state radiation is a shift in $M(nj)$ towards lower values, as evident in Fig. 10 (C), which was taken into account in the analysis by setting a loose $M(nj)$ interval cut. An event was accepted if its jets invariant mass was in the range $0.55\,M_{S}\leq M(nj)\leq 1.05\,M_{S}$ for the $ZZ,WW$ channels and $0.8\,M_{S}\leq M(nj)\leq 1.05\,M_{S}$ for the $hh$ channel. Based on Fig. 10 (C), the $hh$ channel accepted range was adopted to be narrower as compared to that of the $ZZ,WW$ channels. Most likely, this is because jets coming from a Higgs boson’s decay are produced more collinearly, and hence lose less energy in the form of radiation. This fact is supported by examining in Fig. 10 (A), that shows the tendency of jet pairs coming from the Higgs boson’s decay to accumulate towards lower $P_{T}$ values in comparison to those coming from $WW,ZZ$.
@@ 4.4 $2l2q$ Final States
The $2l2q$ channel is also exclusively produced by $S\rightarrow ZZ$. All possible backgrounds were generated in MG5_aMC via $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}l^{+}l^{-}q\overline{q}$. As in the previous channels, the most crucial distinction between signal and background is the invariant mass which separates the resonant production from the background.
@@
On the leptonic side, we required $P_{T}(l^{\pm})\geq\frac{1}{5}M_{Z}$, based on a direct comparison of the transverse momentum of single leptons, to guarantee that we did not lose any possible leptons resulting from $Z$ decays. Then, we paired the same-flavour oppositely-charged leptons if their invariant mass satisfied $0.75M_{Z}\leq M(l^{+}l^{-})\leq 1.25M_{Z}$. Similarly, on the jet side, we asked for $0.4M_{Z}\leq M(nJ)\leq 1.5M_{Z}$, and then paired jets according to their transverse momentum $P_{T}(J)\geq\frac{1}{5}M_{Z}$, which is also confirmed from single $P_{T}(J)$ comparison between signal and background. We then look for resonant production around the scalar mass by computing the invariant mass of all possible combinations of the accepted ($l^{+}l^{-}$) pairs, together with all the accompanying jets ($nJ$) in each event using the formula,
@@
such that $P(J)=\sum_{k}p(J_{k})$ is the previously defined four vectors that add up all the jets present in a single event. This approach returned the correct invariant mass, which shifted towards lower mass values because of the unavoidable final-state radiation, especially towards higher $M_{s}$ values, as depicted in Fig. 12 (C). The same figure also suggests a selection range of $M_{s}-150\,~{}\mathrm{GeV}\leq M(l^{+}l^{-}+nJ)\leq M_{s}+10\,~{}\mathrm{GeV}$. The event is then selected if it contains at least one combination that passe the $M(l^{+}l^{-}+nJ)$ cut and has at least one $l^{+}l^{-}$ pair and one ($J\overline{J}$) pair passing the previous $M_{Z}(l^{+}l^{-},nJ),P_{T}(l^{+}l^{-},J\overline{J})$ constraints. Following the invariant mass cuts, the remaining events were subjected to cuts on $P_{T}(l^{+}l^{-},J\overline{J}),\,\eta(l^{+}l^{-},J\overline{J})$, through direct comparison between the signal and background for each specific $M_{s}$ value, as shown in Fig. 12 (A,B) for $M_{s}=500$ GeV. The $P_{T}$ bins for this channel were selected by requiring $N_{s}\geq 2N_{b}$, and $N_{s}\geq 1.8N_{b}$ for the pseudorapidity ($\eta$) bins.
@@
Fluctuations in $P_{T},\eta$ were then considered according to Eq. 36, by repeating the previous analysis for different invariant mass ranges $M_{s}-250$ GeV $\leq M(l^{+}l^{-}+nJ)\leq M_{s}+20$ GeV which led to the sensitivity curve for this channel, plotted in Fig. 13 at $L=10^{4}$ fb-1.
@@ 4.5 $2\gamma 2q$ Final States
This final state arises from the $S\rightarrow h_{1}h_{1}$ decay channel in addition to all the different, non resonant background channels that lead to the same final state at the muon collider ($\mu^{+}\mu^{-}\longrightarrow\mu^{+}\mu^{-}2\gamma 2J$), which was generated in MG5_aMC. This channel is very challenging because of the difficulty in reconstructing both photons and jets in the final state because a considerable number of photons would be present from the muon and quark QED radiation.
@@
As before, the main difference between the signal and background is the absence of resonant production in the background, which suggests invariant mass cuts on the final state as an effective probe to identify the signal. Photons resulting from Higgs boson decay could be separated from photons resulting from other decay modes, or initial/final state radiation, by requiring $P_{T}(\gamma)\geq\frac{2}{5}M_{h}$, which was confirmed through direct comparison of the transverse momentum distributions. Then, in order to pair photons we computed the invariant mass of all the possible combinations of the photons that passed the $P_{T}$ cut, and only keeping the pairs that had $0.8M_{h}\leq M(\gamma\gamma)\leq 1.2M_{h}$. We then computed the invariant mass of the selected photon pairs and all jets contained in the event, $M(\gamma\gamma+nJ)$, given by [^6] This is a modification of Eq. (38), where we take into account all the possible combinations of the accepted $\gamma\gamma$ pairs with the entire $n$-jets available in the event.
@@
The event was accepted if it possessed $M_{s}-80\,~{}\mathrm{GeV}\leq M(\gamma\gamma+nJ)\leq M_{s}+80\,~{}\mathrm{GeV}$, and contained at least one accepted $\gamma\gamma$ pair. Subsequently, jet pairs were constructed for all possible combinations of the remaining jets after the $P_{T}(J)\geq\frac{2}{5}$ constraint was applied.
@@
The accepted events were further subjected to transverse momentum and pseudorapidity constraints, obtained through a direct comparison between the remaining signal and background events, as shown in Fig. 40 (A,B), for $M_{s}=600$ GeV. We automated these cuts to make them sensitive to the $M_{s}$ value, as shown in Table 6. The $P_{T}$ bins were obtained by requiring $N_{s}\geq 1.1N_{b}$, while $N_{s}\geq 1.01N_{b}$ was adopted for pseudorapidity ($\eta$). These observables are, in turn, dependent on the invariant mass, $M(\gamma\gamma+nJ)$, cut applied, so we repeated the above analysis for a different range, $M_{s}-150\,\mathrm{GeV}\leq M(\gamma\gamma+nJ)\leq M_{s}+100\,\mathrm{GeV}$, and computed the resulting uncertainty in the signal and background efficiencies ($\varepsilon_{s},\varepsilon_{b}$) using Eq. (36). This set of cuts significantly suppressed the background, where the background efficiency was $\varepsilon_{b}\sim 1\%$, against the signal efficiency of $\varepsilon_{s}\sim 50\%$ on average. These results were reflected in the characteristic sensitivity curve for the $2\gamma 2q$-channel as plotted in Fig. 15 at a luminosity of $10^{4}$ fb-1.
@@ 4.6 $2l2\nu_{l}$ Final states
The $2l2\nu_{l}$ final state can originate either from the $S$ decay into weak gauge bosons $S\rightarrow ZZ,\,W^{+}W^{-}$, or other backgrounds that can be generated at a muon collider through $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}(\nu_{\mu}\overline{\nu}_{\mu})2l2\nu_{l}$, and all were considered in our analysis. The invariant mass of such final states is not immediately calculable, because of the missing energy carried away by neutrinos. However, unlike in hadron colliders, the partonic energy of the muon beam used in scalar production can be precisely estimated, giving us an opportunity to reconstruct the parent particle’s mass, emphasizing one of the advantages of muon colliders over hadron colliders.
@@
where $M(l^{+}l^{-})$ is the invariant mass of the two oppositely charged leptons. In the case of ZZF, the $l^{+}l^{-}$ pair will originate from a single $Z$-boson, so only pairs with $0.8M_{Z}\leq M(l^{+}l^{-})\leq 1.2M_{Z}$ were accepted. For WWF, $l^{+}$ and $l^{-}$ will originate from different particles, that is,  $W^{+},W^{-}$ respectively, and hence $M(l^{+}l^{-})$ is not expected to peak around the $M_{W}$ mass. Despite this, we found that the signal from WWF tends to outnumber the background in the region $250\,~{}\mathrm{GeV}\leq M(l^{+}l^{-})\leq 600\,~{}\mathrm{GeV}$. In both cases, ZZF and WWF, only leptons with $P_{T}(l^{\pm})\geq\frac{1}{5}M_{X}$ were included in pair formation. In this way, it is most likely that leptons included are those resulting from mediator ($X$) decay. Nevertheless, some percentages of forward muons may satisfy these requirements. Therefore, a peak is expected to occur at $m_{ij}\sim\sqrt{s}$. However, this is not of interest to us, since we are already limiting $M_{S}\in[200,1000]$ GeV, and consequently such peaks will automatically be cut out (see Fig. 16 (C)). We accepted events that had $0.6M_{X}\leq M(l^{+}l^{-}+\Delta)\leq 1.2M_{X}$ due to the broadening of the resulting invariant mass distributions observed in Fig. 16 (C).
@@
The accepted events up to this point were further constrained by applying cuts on $P_{T}(l^{+}l^{-}),\eta(l^{+}l^{-})$, from a signal-background comparison, automated for each $M_{s}$ value as shown in Fig. 16 (A,B) for $M_{s}=500$ GeV. The results are shown in Table 7 for a sample of different $M_{s}$ values. The $P_{T}$ bins were selected by requiring $N_{s}\geq 1.2N_{b}$ and $N_{s}\geq 1.1N_{b}$ for the $\eta$ bins for both the $ZZ,WW$ channels. Similar to the previous channels, the uncertainty in $\varepsilon_{s},\varepsilon_{b}$ was computed from Eq. (36), by repeating the previous analysis for different invariant mass range, $0.5M_{X}\leq M(l^{+}l^{-}+\Delta)\leq 1.3M_{X}$.
@@
The overall analysis yielded a background efficiency of $\varepsilon_{b}\sim 4\%$ against a signal efficiency of $\varepsilon_{s}\sim 40\%$ for the $ZZ\rightarrow 2l2\nu_{l}$. Instead, for the $WW\rightarrow 2l2\nu_{l}$ final state, the results were not as good, because the leptons did not originate from the same particle, which led to less efficient signal-background separation. This is reflected in the sensitivity plots of Fig. 18, which show a more uncertain curve for the $WW$ channel in comparison to the $ZZ$ channel.

### v2

@@ 4.2 $4l$ Final States
The four-lepton final state originates from $S\rightarrow ZZ$, and its dominant background is $\mu^{+}\mu^{-}\rightarrow 2l^{+}2l^{-}\mu^{+}\mu^{-}$ and $\mu^{+}\mu^{-}\rightarrow 2l^{+}2l^{-}\nu_{\mu}\bar{\nu}_{\mu}$. The signal leptons are expected to have a higher transverse momentum compared to the background ones, as they originate from $Z$-boson decays, which suggests a constraint on the transverse momentum in terms of $Z$-boson mass as $\sim P_{T}(l^{\pm})\geq\frac{2}{5}M_{z}$. This assumption was verified by comparing the signal-background transverse momentum distributions for the charged leptons, which showed signal domination in the region $\sim 50\,\mathrm{GeV}\leq P_{T}(l^{\pm})\leq 400\,\mathrm{GeV}$. Consequently, we identified any oppositely charged same-flavour leptons whose $P_{T}(l^{\pm})\geq\frac{2}{5}M_{z}$, and their invariant mass peaked around the $Z$-boson mass, $0.8M_{Z}\leq M(l^{+}l^{-})\leq 1.2M_{Z}$ as valid pairs resulting from $Z-$boson decay. This works as a strong signal-background separator because the leptons produced from the background generated in MG5_aMC@NLO are sourced from different mediators, as discussed in the previous section. This is in contrast to the signal case where they are mainly produced from $Z$-bosons, $S\rightarrow ZZ$. Then, the surviving pairs were further constrained by requiring the invariant mass of the two pairs to be in the vicinity of the scalar mass, $M_{s}-50\,\mathrm{GeV}\leq M(l^{+}l^{-}l^{+}l^{-})\leq M_{s}+50\,\mathrm{GeV}$, which further suppressed the background events, as they were not resonantly produced in this range.
@@
Finally, the resulting sensitivity plot for a luminosity of $10^{4}~{}\mathrm{fb}^{-1}$ is shown in Fig. 9, at $68\%$ and $95\%$ C.L.
@@ 4.3 $4q$ Final States
The main background is similar to those discussed in the previous channel, with $l\rightarrow q$. In principle, the jets originating from ($ZZ,\,\,W^{+}W^{-},\,\,hh$) decays will have almost identical kinematics due to the approximate symmetry in the $Z,W^{\pm},h$ masses especially for $M_{S}\gg M_{Z},M_{W^{\pm}},M_{h}$. Nevertheless, the differences in the branching ratios of Higgs boson and weak gauge boson decays into quarks will lead to a deviation in the kinematical behavior of the two cases, especially for the transverse momentum and invariant mass distributions, as can be seen in Fig. 10 (A,C). These differences can be further investigated by tracking the jet’s origin, and this can play a vital role in exploring the $\mathbb{Z}_{2}$ nature of xSM. Hence, this specific channel, which has not been explored sufficiently to date, to the best of our knowledge, can further constrain the parameter space of FOEWPT. This proves to be an additional advantage of a muon collider, in which such a channel could be explored with much less effort than at hadron colliders due to the lower QCD backgrounds. A detailed study of these differences is left for future work. Jets were clustered using FastJet (v3.3.2) Cacciari:2011ma , where the anti-kT algorithm Catani:1993hr ; Ellis:1993tq ; Cacciari:2008gp with a radius parameter $R=0.4$ was chosen to be the default jet clustering algorithm. The calculation of the invariant mass for this channel is more challenging than in the previous one, mainly due to the high tendency of quarks to radiate via QCD. This means that we cannot simply calculate the invariant mass as previously using ${M^{2}(4j)=\big{(}\sum_{i=1}^{4}p(J_{i})\big{)}^{2}}$. Instead, for each event, we defined ${P=\sum_{i=1}^{n}p(J_{i})}$, that sums all the jets to a single four vector, from which we construct $M^{2}(nj)=P^{2}$. This, still, will not add up to the parent $S$ particle mass, as some of the energy will have escaped in the form of radiation, but this could be neglected as a similar effect will also be found in background events. The main effect of this final-state radiation is a shift in $M(nj)$ towards lower values, as evident in Fig. 10 (C), which was taken into account in the analysis by setting a loose $M(nj)$ interval cut. An event was accepted if its jets invariant mass was in the range $0.55\,M_{S}\leq M(nj)\leq 1.05\,M_{S}$ for the $ZZ,WW$ channels and $0.8\,M_{S}\leq M(nj)\leq 1.05\,M_{S}$ for the $hh$ channel. Based on Fig. 10 (C), the $hh$ channel accepted range was adopted to be narrower compared to that of the $ZZ,WW$ channels. Most likely, this is because jets coming from a Higgs boson’s decay are produced more collinearly, and hence lose less energy in the form of radiation. This fact is supported by examining in Fig. 10 (A), that shows the tendency of jet pairs coming from the Higgs boson’s decay to accumulate towards lower $P_{T}$ values in comparison to those coming from $WW,ZZ$.
@@ 4.4 $2l2q$ Final States
The $2l2q$ channel is also exclusively produced by $S\rightarrow ZZ$. All possible backgrounds were generated in MG5_aMC via $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}l^{+}l^{-}q\overline{q}$. As in the previous channels, the most crucial distinction between signal and background is the invariant mass, which separates the resonant production from the background.
@@
On the leptonic side, we required $P_{T}(l^{\pm})\geq\frac{1}{5}M_{Z}$, based on a direct comparison of the transverse momentum of single leptons, to guarantee that we did not lose any possible leptons resulting from $Z$ decays. Then, we paired the same-flavour oppositely charged leptons if their invariant mass satisfied $0.75M_{Z}\leq M(l^{+}l^{-})\leq 1.25M_{Z}$. Similarly, on the jet side, we asked for $0.4M_{Z}\leq M(nJ)\leq 1.5M_{Z}$, and then paired jets according to their transverse momentum $P_{T}(J)\geq\frac{1}{5}M_{Z}$, which is also confirmed from single $P_{T}(J)$ comparison between signal and background. We then look for resonant production around the scalar mass by computing the invariant mass of all possible combinations of the accepted ($l^{+}l^{-}$) pairs, together with all the accompanying jets ($nJ$) in each event using the formula,
@@
such that $P(J)=\sum_{k}p(J_{k})$ is the previously defined four vectors that add up all the jets present in a single event. This approach returned the correct invariant mass, which shifted towards lower mass values because of the unavoidable final-state radiation, especially toward higher $M_{s}$ values, as depicted in Fig. 12 (C). The same figure also suggests a selection range of $M_{s}-150\,~{}\mathrm{GeV}\leq M(l^{+}l^{-}+nJ)\leq M_{s}+10\,~{}\mathrm{GeV}$. The event is then selected if it contains at least one combination that passes the $M(l^{+}l^{-}+nJ)$ cut and has at least one $l^{+}l^{-}$ pair and one ($J\overline{J}$) pair passing the previous $M_{Z}(l^{+}l^{-},nJ),P_{T}(l^{+}l^{-},J\overline{J})$ constraints. Following the invariant mass cuts, the remaining events were subjected to cuts on $P_{T}(l^{+}l^{-},J\overline{J}),\,\eta(l^{+}l^{-},J\overline{J})$, through direct comparison between the signal and background for each specific $M_{s}$ value, as shown in Fig. 12 (A,B) for $M_{s}=500$ GeV. The $P_{T}$ bins for this channel were selected by requiring $N_{s}\geq 2N_{b}$, and $N_{s}\geq 1.8N_{b}$ for the pseudorapidity ($\eta$) bins.
@@
The fluctuations in $P_{T},\eta$ were then considered according to Eq. 36, by repeating the previous analysis for different invariant mass ranges $M_{s}-250$ GeV $\leq M(l^{+}l^{-}+nJ)\leq M_{s}+20$ GeV which led to the sensitivity curve for this channel, plotted in Fig. 13 at $L=10^{4}$ fb-1.
@@ 4.5 $2\gamma 2q$ Final States
This final state arises from the $S\rightarrow h_{1}h_{1}$ decay channel in addition to all the different, nonresonant background channels that lead to the same final state at the muon collider ($\mu^{+}\mu^{-}\longrightarrow\mu^{+}\mu^{-}2\gamma 2J$), which was generated in MG5_aMC. This channel is very challenging due to the difficulty in reconstructing both photons and jets in the final state, since a considerable number of photons would be present from the muon and quark QED radiation.
@@
As before, the main difference between the signal and background is the absence of resonant production in the background, which suggests invariant mass cuts on the final state as an effective probe to identify the signal. Photons resulting from the Higgs boson decay could be separated from photons resulting from other decay modes, or initial/final-state radiation, by requiring $P_{T}(\gamma)\geq\frac{2}{5}M_{h}$, which was confirmed through direct comparison of the transverse momentum distributions. Then, to pair photons, we computed the invariant mass of all the possible combinations of the photons that passed the $P_{T}$ cut, and only kept the pairs that had $0.8M_{h}\leq M(\gamma\gamma)\leq 1.2M_{h}$. We then computed the invariant mass of the selected photon pairs and all jets contained in the event, $M(\gamma\gamma+nJ)$, given by, [^6] This is a modification of Eq. (38), where we take into account all the possible combinations of the accepted $\gamma\gamma$ pairs with the entire $n$-jets available in the event.
@@
The event was accepted if it had $M_{s}-80\,~{}\mathrm{GeV}\leq M(\gamma\gamma+nJ)\leq M_{s}+80\,~{}\mathrm{GeV}$, and contained at least one accepted $\gamma\gamma$ pair. Subsequently, jet pairs were constructed for all possible combinations of the remaining jets after the $P_{T}(J)\geq\frac{2}{5}$ constraint was applied.
@@
The accepted events were further subjected to transverse momentum and pseudorapidity constraints, obtained through a direct comparison between the remaining signal and background events, as shown in Fig. 14 (A,B), for $M_{s}=600$ GeV. We automated these cuts to make them sensitive to the $M_{s}$ value, as shown in Table 6. The $P_{T}$ bins were obtained by requiring $N_{s}\geq 1.1N_{b}$, while $N_{s}\geq 1.01N_{b}$ was adopted for pseudorapidity ($\eta$). These observables are, in turn, dependent on the invariant mass, $M(\gamma\gamma+nJ)$, cut applied, so we repeated the above analysis for a different range, $M_{s}-150\,\mathrm{GeV}\leq M(\gamma\gamma+nJ)\leq M_{s}+100\,\mathrm{GeV}$, and computed the resulting uncertainty in the signal and background efficiencies ($\varepsilon_{s},\varepsilon_{b}$) using Eq. (36). This set of cuts significantly suppressed the background, where the background efficiency was $\varepsilon_{b}\sim 1\%$, compared to the signal efficiency of $\varepsilon_{s}\sim 50\%$ on average. These results were reflected in the characteristic sensitivity curve for the $2\gamma 2q$-channel as plotted in Fig. 15 at a luminosity of $10^{4}$ fb-1.
@@ 4.6 $2l2\nu_{l}$ Final states
The $2l2\nu_{l}$ final state can originate either from the $S$ decay into weak gauge bosons $S\rightarrow ZZ,\,W^{+}W^{-}$, or other backgrounds that can be generated at a muon collider through $\mu^{+}\mu^{-}\rightarrow\mu^{+}\mu^{-}(\nu_{\mu}\overline{\nu}_{\mu})2l2\nu_{l}$, and all were considered in our analysis. The invariant mass of such final states is not immediately calculable, because of the missing energy carried away by neutrinos. However, unlike in hadron colliders, the energy of the muon beam used in scalar production can be precisely estimated, giving us an opportunity to reconstruct the parent particle’s mass, emphasizing one of the advantages of muon colliders over hadron colliders.
@@
where $M(l^{+}l^{-})$ is the invariant mass of the two oppositely charged leptons. In the case of ZZF, the $l^{+}l^{-}$ pair will originate from a single $Z$-boson, so only pairs with $0.8M_{Z}\leq M(l^{+}l^{-})\leq 1.2M_{Z}$ were accepted. For WWF, $l^{+}$ and $l^{-}$ will originate from different particles, that is, $W^{+},W^{-}$ respectively, and therefore $M(l^{+}l^{-})$ is not expected to peak around the $M_{W}$ mass. Despite this, we found that the signal from WWF tends to outnumber the background in the region $250\,~{}\mathrm{GeV}\leq M(l^{+}l^{-})\leq 600\,~{}\mathrm{GeV}$. In both cases, ZZF and WWF, only leptons with $P_{T}(l^{\pm})\geq\frac{1}{5}M_{X}$ were included in pair formation. In this way, it is most likely that leptons included are those resulting from mediator ($X$) decay. However, a fraction of forward muons may satisfy these requirements. Therefore, a peak is expected to occur at $m_{ij}\sim\sqrt{s}$. However, this is not of interest to us, since we are already limiting $M_{S}\in[200,1000]$ GeV, and consequently such peaks will automatically be cut out (see Fig. 16 (C)). We accepted events that had $0.6M_{X}\leq M(l^{+}l^{-}+\Delta)\leq 1.2M_{X}$ due to the broadening of the resulting invariant mass distributions observed in Fig. 16 (C).
@@
The accepted events up to this point were further constrained by applying cuts on $P_{T}(l^{+}l^{-}),\eta(l^{+}l^{-})$, from a signal-background comparison, automated for each $M_{s}$ value as shown in Fig. 16 (A,B) for $M_{s}=500$ GeV. The results are shown in Table 7 for a sample of different $M_{s}$ values. The $P_{T}$ bins were selected by requiring $N_{s}\geq 1.2N_{b}$ and $N_{s}\geq 1.1N_{b}$ for the $\eta$ bins for the $ZZ,WW$ channels. Similarly to the previous channels, the uncertainty in $\varepsilon_{s},\varepsilon_{b}$ was computed from Eq. (36), by repeating the previous analysis for different invariant mass range, $0.5M_{X}\leq M(l^{+}l^{-}+\Delta)\leq 1.3M_{X}$.
@@
The full analysis yielded a background efficiency of $\varepsilon_{b}\sim 4\%$ against a signal efficiency of $\varepsilon_{s}\sim 40\%$ for $ZZ\rightarrow 2l2\nu_{l}$. Instead, for the $WW\rightarrow 2l2\nu_{l}$ final state, the results were not as good, because the leptons did not originate from the same particle, which led to a less efficient signal-background separation. This is reflected in the sensitivity plots in Fig. 18, which show a more uncertain curve for the $WW$ channel compared to the $ZZ$ channel.